- Recording to HOB+0035. +3 GB
- Usual RFI from the NBN in last two channels
- Recording to HOB+0122. +20792 GB
- RFI ongoing in S band.
- Recording to HOB+0120. +22656 GB
- Antenna was stuck for about an hour before the experiment. Onsite staff managed to return control about 15 min into exp. First source was 244-0716b.
- Recording to USN-0027. +3 GB
- Receiver is warm, so tsys higher than usual.
- Status monitor clock shows wrong datetime. However, log time agrees with monit2 agrees with UT. So will clarify inconsistency at a later time.